OpenStudy (anonymous):

factor out the greatest common factor of 8x^2-12x^3

OpenStudy (anonymous):

When you factor you have to remember the first most important rule: Look for the greatest common number and/or variable. you have to find the greatest common factor of 8 and 12 and x^2 and x 8 and 12 gcf comes out to be 4 and x^2 and x comes out to be x Then you have to take 4x out of the whole expression 4x (2x + 3) If you were to multiply this out by foil, you should get your original expression :)

OpenStudy (anonymous):

What if you want to factor out a negative

OpenStudy (anonymous):

8X^2-12X^3=4(2x^2-3x^3)=4X^2(2-3x) So 4x^2 is the greatest factor

OpenStudy (anonymous):

then see what 2 factors get you to the negative